Solve -y''-2y'+2y = e-2x

y(0)=1

y(infinity )=0


1
Expert's answer
2021-04-15T07:41:01-0400

"-y''-2y'+2y=e^{-2x}"

"y(0)=1"

"y(\infty)=0"

Solution:

"y''+2y'-2y=-e^{-2x}"

Characteristic equation:

"r^2+2r-2=0" ;

"D=\sqrt{4+4\cdot 2}=2\sqrt3" ;

"r_1=-1+\sqrt3" ; "r_2=-1-\sqrt3" .

The general solution of the homogeneous differential equation:

"y_0=C_1e^{(-1+\sqrt3)x}+C_2e^{(-1-\sqrt3)x}"

Let's find a partial solution of the non-homogeneous equation in the form

"y_p=Ae^{-2x}"

"y_p'=-2Ae^{-2x}"

"y_p''=4Ae^{-2x}"

Substitute "y_p", "y_p'", "y_p''" into the origin differential equation:

"4Ae^{-2x}-4Ae^{-2x}-2Ae^{-2x}=-e^{-2x}"

"A=\frac12"

The general solution of the non-homogeneous differential equation:

"y=y_0+y_p=C_1e^{(-1+\sqrt3)x}+C_2e^{(-1-\sqrt3)x}+\frac12e^{-2x}"

Let's apply conditions:

"y(0)=C_1+C_2+\frac12=1"

"y(\infty)=0\implies" "C_1=0" (because if "C_1\neq0" then "y(\infty)\to\infty" )

"C_2=\frac12"

Answer: "y=\frac12e^{(-1-\sqrt3)x}+\frac12e^{-2x}" .